OpenStudy (anonymous):

Serena parachutes from a plane at a height of 20,000 ft above sea level. Several minutes later she lands on the ground 15 ft below sea level . What was the total vertical distance that Serena parachuted? @ganeshie8

ganeshie8 (ganeshie8):

you started at a height of `20,000 ft ` and landed at a height of `15 ft` so how much is the chang ein height ?

OpenStudy (anonymous):

19985?

ganeshie8 (ganeshie8):

Perfect ! you simply subtract : change in height = \(\large 20000-15 = 19985\)

OpenStudy (anonymous):

how do i write i thoe

ganeshie8 (ganeshie8):

write what

OpenStudy (anonymous):

the number

ganeshie8 (ganeshie8):

not getting u

OpenStudy (anonymous):

do i write it like this 19,985

OpenStudy (anonymous):

The variable t represents the temperature at 6 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Washington dropped 19°F between 6 p.m. and 10 p.m. The temperature at 10 p.m. was –12°F. What was the temperature at 6 p.m.? A. 12 – 19 = t B. 19 + 12 = t C. t – 19 = –12 D. 19 – t = –12

OpenStudy (anonymous):

can you check it i think its b

OpenStudy (anonymous):

@ganeshie8